OpenStudy (anonymous):

given that g(x)=x^2-3x-6, find (g of g)(-3)

OpenStudy (anonymous):

First compute g(3), then compute g(of that #) g(3) is the value of g when x = -3, so g(-3) = (-3)^2 -3(-3) - 6 = 9+9-6 = 12 Now, you find g(12).

OpenStudy (anonymous):

thank you i knew to plug in the -3 value but i wasn't sure what to do after i got the first answer

OpenStudy (anonymous):

welcome.